OpenStudy (anonymous):

According to records, the amount of precipitation in a certain city on a November day has a mean of .1 inches, with a standard deviation of .6 inches. What is the probability that the mean daily precipitation will be .098 inches or more for a random sample of 40 November days (taken over many years)?

OpenStudy (debbieg):

This is asking about a probability related to a SAMPLE MEAN, not just a data value. Thuse, you need to use the Central Limit Theorem formula to get the z-score: \[\large z=\frac{ \bar{x} -\mu}{ \frac{ \sigma }{ \sqrt{n}} }\] You have \[\bar{x}=0.098\]\[\mu=0.1\]\[\sigma=0.6\]and n = 40. Compute the z-score, then use a normal table (or a calculator) to find the area to the right of the z-score.

OpenStudy (anonymous):

i get an answer of 0.021

OpenStudy (anonymous):

wait thats just the z score right?

OpenStudy (debbieg):

Yes. And you should have gotten z=-0.021, since x-bar<mu. So now you need the standard normal probability for z>-0.021.

OpenStudy (anonymous):

so on the calculator i would just do P(Z<=-.021) correct?

OpenStudy (anonymous):

or would it be 1 - that?

OpenStudy (debbieg):

What is the probability that the mean daily precipitation will be .098 inches OR MORE... so you want the P(Z>-0.021) ---- correct, you can get the area to the left, and subtract from 1. Or get the area to the right.

OpenStudy (debbieg):

Or, if you understand how to take advantage of the symmetry of the distribution, you can even use P(Z<0.021) which is = P(Z>-0.021). :)

OpenStudy (anonymous):

awesome I think i am following you